Consider the following table: GDP per capita (constant dollars) India USA 470 1985 32,300 1995 38,680 650 2005 48,760 1,010 2015 51,490 1,810 1. (9 points) Fill in the table below. Annualized growth rate India USA 1985 to 1995 1995 to 2005 2005 to 2015 (3 points) Output per person is close to output per hour. What do we call output per hour? 2.
3. (3 points) What are the three components that determine long run growth? (2 points) Which of these three components likely explain the differential growth rates 4. between USA and India over the past 30 years? (4 points) Suppose the two nations embrace strong growth as their goal and they make 5. intelligent decisions over the next 20 years. Which of the two nations is likely to witness the slowest GDP/capita growth and which is likely to achieve the fastest GDP/capita growth? Briefly explain why.
Answer #1

1. Annualized Growth rate

USA India
1985-1995 38680 - 32300 / 32300 * 100 = 19.75% 650-470/470 * 100 = 38.30%
1995-2005 48760 - 38680 / 38680 * 100 = 26.06% 1010 - 650 / 650 * 100 = 55.38%
2005-2015 51490 - 48760 / 48760 * 100 = 5.60% 1810 - 1010 / 1010 *100 = 79.21%

2. Output per hour is the output delivered by certain set of people in an hour.

3. The three components that determine long run growth are :

Technological Advancement

Population increase

Increase in market size

4. Differential growth rates between USA and India are :

Technological advancement level is high in USA than India.

Government support for growth in GDP

Infrastructure development is high in USA

5. India will achieve the fastest GDP growth as they have very high scope to improve the GDP whereas, the scope is very less in USA to improve the GDP growth further as already it is very high.